# A novel multipoint stress control volume method for linear elasticity on quadrilateral grids

### v1

@@ Abstract
In this paper, we develop a novel control volume method that is locally conservative and locking-free for linear elasticity problem on quadrilateral grids. The symmetry of stress is weakly imposed through the introduction of a Lagrange multiplier. As such, the method involves three unknowns: stress, displacement and rotation. To ensure the well-posedness of the scheme, a pair of carefully defined finite element spaces is used for the stress, displacement and rotation such that the inf-sup condition holds. An appealing feature of the method is that piecewise constant functions are used for the approximations of stress, displacement and rotation, which greatly simplifies the implementation. In particular, the stress space is defined delicately such that the stress bilinear form is localized around each vertex, which allows for the local elimination of the stress, resulting in a cell-centered system. By choosing different definitions of the space for rotation, we develop two variants of the method. In particular, the first method uses a constant function for rotation over the interaction region, which allows for further elimination and results in a cell-centered system involving displacement only. A rigorous error analysis is performed for the proposed scheme. We show a quasi-optimal approximation (up to data-oscillation term) for $L^{2}$-error of the stress and rotation without requiring additional regularity assumption. Moreover, we can also prove the superconvergence for discrete $H^{1}$-error and $L^{2}$-error of displacement. Extensive numerical simulations indicate that our method is efficient and accurate, and can handle problems with discontinuous coefficients.
@@ 1 Introduction
In this paper we aim to develop a new method that only uses piecewise constant approximations for the involved unknowns, that can be further reduced to a symmetric positive definite system. Our method is closely related to the multipoint stress mixed finite element method proposed in [4, 5], but with much simpler construction and implementation in the sense that no special quadrature rule is needed. The devising of the stress space is more subtle as it needs to be carefully balanced with the displacement such that the inf-sup condition holds. To this end, we divide the quadrilateral element into four smaller quadrilaterals by connecting the interior points to the midpoint of each edge and then define the stress space as a constant function over each smaller quadrilateral, and at the same time it is normal continuous over the edges lying on the quadrilaterals, but no continuity is imposed for the new edges generated by the subdivision. As a consequence, the bilinear form associated with the stress is localized around each vertex, which resembles the vertex quadrature rule used in multipoint stress mixed finite element method. The major difference is that we use piecewise constant function to avoid the quadrature rule. Then we develop two variants of the method by choosing different spaces for the rotation. In the first method, we let the rotation be a constant function over each interaction region formed by the four smaller quadrilaterals sharing the original vertex. Then the system can be further reduced to achieve a symmetric and positive definite cell-centered system for the displacement only. In the second method, we choose the rotation to be a constant function over each quadrilateral. In this case, the stress can not be further eliminated, but this variant can handle discontinuous coefficients accurately. In particular, a simple element-wise reconstruction will lead to $H(\mathrm{div};\Omega)$-conforming stress. Our method can also be viewed as a mixed-type finite volume method, but we can carry out the error analysis in the framework of finite element methods. This can simplify the analysis as it is well-known it is not always easy to analyze finite volume method.
@@
We perform a rigorous error analysis for the proposed scheme. The unique solvability of the solution is established. Moreover, we also show the uniform boundedness of the solution, with respect to $\lambda$. In many practical scenarios, the solution may have limited regularity, however, standard error analysis requires the solution to be smooth enough. To this end, we prove the convergence of $L^{2}$-error of the stress and rotation, and the discrete $H^{1}$-error of displacement without requiring additional regularity. We emphasize that we do not invoke the medius analysis [16, 15] to show the convergence under minimal regularity assumption, instead, we use the moment-preserving operator to build a link between the discrete formulation and the weak formulation. Our method also inherits the locking-free property of standard mixed finite element method, which is also analyzed rigorously. Besides, we also prove the superconvergence of $L^{2}$-error of displacement under the assumption that the solution is smooth enough. In summary, our method owns several appealing features, which include: (1) the method enjoys local conservation, and a simple reconstruction will yield $H(\mathrm{div};\Omega)$-conforming stress; (2) it is locking-free, which allows to handle problems with nearly incompressible materials; (3) piecewise constant functions are used for all the unknowns, which makes the implementation easy; (4) the stress bilinear form is block-diagonal, which allows local elimination of stress and reduction to a cell-centered system. All these salient features make our method a good candidate for handling problems arising from practical applications.
@@
    We prove the quasi-optimal approximation (up to data oscillation term) without requiring additional regularity assumption. Moreover, the superconvergence for displacement can be proved under the assumption that the solution is smooth smooth.
@@
The rest of the paper is organized as follows. In the next section, we introduce the model problem and present the corresponding weak formulation. In section 3, we describe the proposed methods. The comprehensive error analysis including the quasi-optimal approximation (up to data-oscillation term) and the superconvergence for $L^{2}$-error of displacement is performed in section 4. In section 6, extensive numerical simulations including three-dimensional tests are carried out to demonstrate the capability, efficiency and superior performance of the proposed methods.
@@ 2 The mathematical model
Here $\lambda$ and $\mu$ are positive, called the Lamé parameters, $\mbox{tr}(\underline{G})$ is the trace of function $\underline{G}$, and $\underline{I}$ is the identity matrix. We define $\|\underline{G}\|_{\mathcal{A}}^{2}=(\mathcal{A}\underline{G},\underline{G})$. Now we introduce some notation that will be used throughout the paper. The norms and seminorms of the Sobolev spaces $W^{k,p}(\mathcal{O}),k\in\mathbb{R},p>0$ are denoted by $\|\cdot\|_{W^{k,p}(\mathcal{O})}$ and $|\cdot|_{W^{k,p}(\mathcal{O})}$, respectively, where $\mathcal{O}\subset\Omega,\Omega\subset\mathbb{R}^{d}$. When $k=1$, we use $H^{r}(\mathcal{O})$ to represent the corresponding space. The spaces of vector- and matrix-valued functions with all the components in $H^{r}(\mathcal{O})$ will be respectively denoted as $\bm{H}^{r}(\mathcal{O})$ and $\underline{H}^{r}(\mathcal{O})$. The corresponding norm and semi-norm are respectively denoted as $\|\cdot\|_{H^{r}(\mathcal{O})}$ and $|\cdot|_{H^{r}(\mathcal{O})}$. We use $(\cdot,\cdot)_{D}$ to represent the standard $L^{2}$-inner product over $D\subset\mathbb{R}^{d}$ and when $D$ coincides with $\Omega$, the subscript will be omitted. The corresponding norm is denoted as $\|\cdot\|_{L^{2}(D)}$. We define $\bm{H}(\mathrm{div};\Omega):=\{\bm{v}\in\bm{L}^{2}(\Omega),\nabla\cdot\bm{v}\in L^{2}(\Omega)\}$, which is equipped with the norm $\|\bm{v}\|_{\mathrm{div}}^{2}=\|\bm{v}\|_{L^{2}(\Omega)}^{2}+\|\nabla\cdot\bm{v}\|_{L^{2}(\Omega)}^{2}$. In addition, we use $\underline{H}(\mathrm{div};\Omega)$ to represent the tensor field where each row belongs to $\bm{H}(\mathrm{div};\Omega)$.
@@ 3.1 Preliminaries
where $\bm{t}_{i}:=(t_{i}^{1},t_{i}^{2})^{T},i=1,\cdots,4$ represents the unit tangential vector of $e_{i}$.
@@ 3.2 The discrete formulation and statement of the main results
Let $(\underline{\sigma},\bm{u})$ be the exact solution of (2.1)-(2.2) and let $(\underline{\sigma}_{h},\bm{u}_{h},\gamma_{h})\in\underline{\Sigma}_{h}\times\bm{U}_{h}\times\Gamma_{h}$ be the discrete solution of (3.2)-(3.4) obtained by either Method 1 or Method 2. Then, there exists a positive constant $C$ independent of the mesh size and $\lambda$ such that
$\|\underline{\sigma}-\underline{\sigma}_{h}\|_{L^{2}(\Omega)}$ $\leq C\Big{(}\inf_{\widehat{\sigma}_{h}\in\underline{\Sigma}_{h}}\|\widehat{\sigma}_{h}-\underline{\sigma}\|_{L^{2}(\Omega)}+\inf_{\widehat{\gamma}_{h}\in\Gamma_{h}}\|\gamma-\widehat{\gamma}_{h}\|_{L^{2}(\Omega)}+\Big{(}\sum_{M\in\mathcal{T}_{M}}h_{M}^{2}\|\bm{f}\|_{L^{2}(M)}^{2}\Big{)}^{\frac{1}{2}}\Big{)},$
$\|(\mathcal{J}_{h}\bm{u}-\bm{u}_{h},\widehat{\gamma}_{h}-\gamma_{h})\|_{h}$ $\leq C\Big{(}\inf_{\widehat{\sigma}_{h}\in\underline{\Sigma}_{h}}\|\widehat{\sigma}_{h}-\underline{\sigma}\|_{L^{2}(\Omega)}+\inf_{\widehat{\gamma}_{h}\in\Gamma_{h}}\|\gamma-\widehat{\gamma}_{h}\|_{L^{2}(\Omega)}+\Big{(}\sum_{M\in\mathcal{T}_{M}}h_{M}^{2}\|\bm{f}\|_{L^{2}(M)}^{2}\Big{)}^{\frac{1}{2}}\Big{)},$
where $\mathcal{J}_{h}$ is defined in (4.2). In addition, if we assume the solution is smooth enough; i.e., $\underline{\sigma}\in\underline{H}^{1}(\Omega)$ and $\mathrm{div}\underline{\sigma}\in\bm{H}^{1}(\Omega)$, then the following superconvergence holds
$\|\bm{Q}_{h}\bm{u}-\bm{u}_{h}\|_{L^{2}(\Omega)}\leq C\Big{(}h^{2}\Big{(}|\underline{\sigma}|_{H^{1}(\Omega)}+\|\mathrm{div}\underline{\sigma}\|_{H^{1}(\Omega)}\Big{)}+h\Big{(}\sum_{M\in\mathcal{T}_{M}}h_{M}^{2}\|\bm{f}\|_{L^{2}(M)}^{2}\Big{)}^{\frac{1}{2}}\Big{)},$
@@ 4 Error analysis
In this section, we show the unique solvability of the discrete formulation (cf. Theorem 4.1) and prove the quasi-optimal approximation for $L^{2}$-error of stress and rotation, and the discrete $H^{1}$-error of displacement (up to data-oscillation term); see Theorem 4.2 and Corollary 4.1.
@@ 4.1 Some elementary properties
$\|\bm{v}\|_{L^{2}(M)}$ and $\|\widehat{\bm{v}}\|_{L^{2}(M)}$ are equivalent uniformly in $h$. Moreover, it is shown in [20, Lemma 5.5] that
@@
We define the following projection operator over each macro-element
$\mathcal{J}_{h}\bm{w}_{|M}=\frac{1}{\bm{q}_{1}h_{e_{1}}+\bm{q}_{2}h_{e_{2}}+\bm{q}_{3}h_{e_{3}}+\bm{q}_{4}h_{e_{4}}}\Big{(}(\bm{w},\bm{q}_{1})_{e_{1}}+(\bm{w},\bm{q}_{2})_{e_{2}}+(\bm{w},\bm{q}_{3})_{e_{3}}+(\bm{w},\bm{q}_{4})_{e_{4}}\Big{)}$ (4.2)
for any $\bm{q}_{i}\in\bm{P}_{0}(e_{i}),e_{i}\in\mathcal{F}_{dl}\cap E,i=1,\cdots,4$. This can be viewed as the weighted average among the dual edges of each macro-element. The following holds
$\sum_{e\in\mathcal{F}_{dl}}(\mathcal{J}_{h}\bm{w},\left\llbracket\underline{\theta}_{h}\bm{n}\right\rrbracket)_{e}=\sum_{e\in\mathcal{F}_{dl}}(\bm{w},\left\llbracket\underline{\theta}_{h}\bm{n}\right\rrbracket)_{e}.$ (4.3)
We can observe that $\mathcal{J}_{h}$ preserves constant over each macro-element, thus an application of the Bramble-Hilbert lemma implies
$\|\bm{\theta}-\mathcal{J}_{h}\bm{\theta}\|_{L^{2}(M)}\leq Ch_{M}\|\nabla\bm{\theta}\|_{L^{2}(M)}\quad\forall\bm{\theta}\in H^{1}(M).$ (4.4)
@@
$(\bm{\Pi}^{u}\bm{v}\cdot\bm{n},\mu)_{e}=(\bm{v}\cdot\bm{n},\mu)_{e}\quad\forall\mu\in P_{0}(e),e\in\mathcal{F}_{pr}^{\frac{1}{2}}\cap\partial E,$ (4.5)
@@
$\|\bm{v}-\bm{\Pi}^{u}\bm{v}\|_{L^{2}(M)}\leq Ch|\bm{v}|_{H^{1}(M)}\quad\forall\bm{v}\in\bm{H}^{1}(M),M\in\mathcal{T}_{M}.$
@@
where $e_{k}\in\mathcal{F}_{pr}^{\frac{1}{2}}$ is the half edge of $e$. In addition, the following holds (cf. [32, (3.9)])
@@
We use $\underline{I}_{h}^{e}$ and $\underline{\Pi}_{h}$ to represent the extension of $\bm{\Pi}^{u}$ and $\bm{\Pi}_{h}$ to the matrix field, i.e., $\underline{I}_{h}^{e}=[\bm{\Pi}^{u}]^{2}$ and $\underline{\Pi}_{h}=[\bm{\Pi}_{h}]^{2}$. We can deduce from the definition of $\underline{\Pi}_{h}$ that
@@
Now we recall the following result from [8].
@@
We define the finite element spaces as follows:
$\bm{V}_{h}:$ =$=\{\bm{q}=(q_{1},q_{2})^{T}\in\bm{H}^{1}(\Omega):(q_{i})_{|E}\leftrightarrow\widehat{q}_{i}\in\mathcal{S}_{2}(\widehat{E}),i=1,2,\forall E\in\mathcal{T}_{h}\},$
$W_{h}:$ =$=\{w\in L^{2}(\Omega):w_{|E}\leftrightarrow\widehat{w}\in Q_{0}(\widehat{E}),\forall E\in\mathcal{T}_{h}\},$
where $\widehat{E}$ is the reference square corresponding to $E$ and
$\mathcal{S}_{2}(\widehat{E})=P_{2}(\widehat{E})+span\{\widehat{x}^{2}\widehat{y},\widehat{x}\widehat{y}^{2}\}.$
Then, there exists a unique solution $(\bm{\rho}_{h},p_{h}^{*})\in\bm{V}_{h}\times W_{h}$ to the following system
$(\nabla\bm{\rho}_{h},\nabla\bm{v})+(p_{h}^{*},\nabla\cdot\bm{v})$ $=(\bm{f},\bm{v})\quad\forall\bm{v}\in\bm{V}_{h},$ (4.13)
$(\mathrm{div}\bm{\rho}_{h},q)$ $=\textnormal{DF}J^{-1}\textnormal{DF}(g,q)\quad\forall q\in W_{h}.$ (4.14)
Moreover, it holds
$\|\bm{\rho}_{h}\|_{H^{1}(\Omega)}\leq C\|g\|_{L^{2}(\Omega)}.$ (4.15)
@@
(inf-sup condition). For both Method 1 and Method 2, there exists a positive constant $C$ independent of the mesh size such that
$\sup_{\underline{0}\neq\underline{w}_{h}\in\underline{\Sigma}_{h}}\frac{\sum_{e\in\mathcal{F}_{pr}^{\frac{1}{2}}}(\underline{w}_{h}\bm{n},\left\llbracket\bm{v}_{h}\right\rrbracket)_{e}+(\text{as}(\underline{w}_{h}),\eta_{h})}{\|\underline{w}_{h}\|_{L^{2}(\Omega)}}\geq C\|(\bm{v}_{h},\eta_{h})\|_{h}\quad\forall(\bm{v}_{h},\eta_{h})\in\bm{U}_{h}\times\Gamma_{h}.$
@@
We apply Piola transformation to the second term, i.e., $(\text{as}(\underline{w}_{h}),\eta_{h})$. Then it holds
$(\text{as}(\underline{w}_{h}),\eta_{h})_{M}=DF(\text{as}(\widehat{w}_{h}),\widehat{\eta}_{h})_{\widehat{M}}.$
We define $\widehat{\underline{\psi}}_{2}$ by
$\widehat{\bm{\psi}}_{2}\cdot\widehat{\bm{n}}_{|e}=\frac{1}{|\widehat{e}|}\int_{\widehat{e}}\mathrm{curl}(\widehat{\bm{\rho}}_{h})\widehat{\bm{n}}\quad\forall e\in\mathcal{F}_{pr}^{\frac{1}{2}}\cap\partial E_{i}.$ (4.18)
Then the use of trapezoidal quadrature rule yields
$(\text{as}(\widehat{\underline{\psi}}_{2}),\widehat{\eta}_{h})_{\widehat{M}}=(\text{as}(\mathrm{curl}(\widehat{\bm{\rho}}_{h})),\widehat{\eta}_{h})_{\widehat{M}}=(\mathrm{div}\widehat{\bm{\rho}}_{h},\widehat{\eta})_{\widehat{M}}=(DF)J^{-1}(\mathrm{div}\bm{\rho}_{h},\eta_{h})_{M},$ (4.19)
where we use the fact that $\mathrm{div}\widehat{\bm{\rho}}_{h}$ restricted to the edge of $\widehat{M}$ is a linear function and the trapezoidal quadrature rule is exact for linear function. Therefore, in view of (4.14) with $g=\eta_{h}-\text{as}(\underline{\psi}_{1})$, we can obtain
@@
Moreover, we have
$\underline{\psi}_{2}\cdot\bm{n}_{|e}=\frac{1}{h_{e}}\widehat{\psi}_{2}\cdot\widehat{\bm{n}}_{|\widehat{e}}.$
This, (4.15), the scaling arguments and (4.17) yield

### v2

@@ Abstract
In this paper, we develop a novel control volume method that is locally conservative and locking-free for linear elasticity problem on quadrilateral grids. The symmetry of stress is weakly imposed through the introduction of a Lagrange multiplier. As such, the method involves three unknowns: stress, displacement and rotation. To ensure the well-posedness of the scheme, a pair of carefully defined finite element spaces is used for the stress, displacement and rotation such that the inf-sup condition holds. An appealing feature of the method is that piecewise constant functions are used for the approximations of stress, displacement and rotation, which greatly simplifies the implementation. In particular, the stress space is defined delicately such that the stress bilinear form is localized around each vertex, which allows for the local elimination of the stress, resulting in a cell-centered system. By choosing different definitions of the space for rotation, we develop two variants of the method. In particular, the first method uses a constant function for rotation over the interaction region, which allows for further elimination and results in a cell-centered system involving displacement only. A rigorous error analysis is performed for the proposed scheme. We show the optimal convergence for $L^{2}$-error of the stress and rotation. Moreover, we can also prove the superconvergence for $L^{2}$-error of displacement. Extensive numerical simulations indicate that our method is efficient and accurate, and can handle problems with discontinuous coefficients.
@@ 1 Introduction
In this paper we aim to develop a new method that only uses piecewise constant approximations for the involved unknowns, that can be further reduced to a symmetric positive definite system. Our method is closely related to the multipoint stress mixed finite element method proposed in [4, 5], but with much simpler construction and implementation in the sense that no special quadrature rule is needed. The devising of the stress space is more subtle as it needs to be carefully balanced with the displacement such that the inf-sup condition holds. To this end, we divide the quadrilateral element into four smaller quadrilaterals by connecting the interior points to the midpoint of each edge and then define the stress space as a constant function over each smaller quadrilateral, and at the same time it is normal continuous over the edges lying on the quadrilaterals, but no continuity is imposed for the new edges generated by the subdivision. As a consequence, the bilinear form associated with the stress is localized around each vertex, which resembles the vertex quadrature rule used in multipoint stress mixed finite element method. The major difference is that we use piecewise constant function to avoid the quadrature rule. Then we develop two variants of the method by choosing different spaces for the rotation. In the first method, we let the rotation be a constant function over each interaction region formed by the four smaller quadrilaterals sharing the original vertex. Then the system can be further reduced to achieve a symmetric and positive definite cell-centered system for the displacement only. In the second method, we choose the rotation to be a constant function over each quadrilateral. In this case, the rotation can not be further eliminated, but this variant can handle discontinuous coefficients accurately. In particular, a simple element-wise reconstruction will lead to $H(\mathrm{div};\Omega)$-conforming stress. Our method can also be viewed as a mixed-type finite volume method, but we can carry out the error analysis in the framework of finite element methods. This can simplify the analysis as it is well-known it is not always easy to analyze finite volume method. We emphasize that the current approach shares a conceptual similarity with the method proposed in [29], particularly in the use of primal and dual grids to ensure method stability. The first method introduced in this paper, after the elimination of stress and rotation variables, results in a smaller system compared to that of [30], potentially leading to faster computations. The second method presented here incorporates displacement and rotation in the final system, with its size determined by the number of primal elements. In contrast, the size of the final system in the method proposed by [30] is dependent on the number of primal edges. Moreover, the method from [30] is capable of handling highly distorted grids, however, the current methods require the $\mathcal{O}(h^{2})$ mesh condition.
@@
We perform a rigorous error analysis for the proposed scheme. The unique solvability of the solution is established. Our method also inherits the locking-free property of standard mixed finite element method. Besides, we also prove the superconvergence of $L^{2}$-error of displacement under the assumption that the solution is smooth enough. In summary, our method owns several appealing features, which include: (1) the method enjoys local conservation, and a simple reconstruction will yield $H(\mathrm{div};\Omega)$-conforming stress; (2) it is locking-free, which allows to handle problems with nearly incompressible materials; (3) piecewise constant functions are used for all the unknowns, which makes the implementation easy; (4) the stress bilinear form is block-diagonal, which allows local elimination of stress and reduction to a cell-centered system. All these salient features make our method a good candidate for handling problems arising from practical applications.
@@
    We prove the convergence error estimates for all the involved variables. Moreover, the superconvergence for displacement can be proved under the assumption that the solution is smooth smooth.
@@
The rest of the paper is organized as follows. In the next section, we introduce the model problem and present the corresponding weak formulation. In section 3, we describe the proposed methods. The comprehensive error analysis and the superconvergence for $L^{2}$-error of displacement are performed in section 4. The extension of the proposed scheme to Darcy equations and the Stokes equations are given in section 5. In section 6, extensive numerical simulations including three-dimensional tests are carried out to demonstrate the capability, efficiency and superior performance of the proposed methods.
@@ 2 The mathematical model
Here $\lambda$ and $\mu$ are positive, called the Lamé parameters, $\mbox{tr}(\underline{G})$ is the trace of function $\underline{G}$, and $\underline{I}$ is the identity matrix. We define $\|\underline{G}\|_{\mathcal{A}}^{2}=(\mathcal{A}\underline{G},\underline{G})$. Now we introduce some notation that will be used throughout the paper. The norms and seminorms of the Sobolev spaces $W^{k,p}(\mathcal{O}),k\in\mathbb{R},p>0$ are denoted by $\|\cdot\|_{W^{k,p}(\mathcal{O})}$ and $|\cdot|_{W^{k,p}(\mathcal{O})}$, respectively, where $\mathcal{O}\subset\Omega,\Omega\subset\mathbb{R}^{d}$. When ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p=2}$, we use $H^{r}(\mathcal{O})$ to represent the corresponding space. The spaces of vector- and matrix-valued functions with all the components in $H^{r}(\mathcal{O})$ will be respectively denoted as $\bm{H}^{r}(\mathcal{O})$ and $\underline{H}^{r}(\mathcal{O})$. The corresponding norm and semi-norm are respectively denoted as $\|\cdot\|_{H^{r}(\mathcal{O})}$ and $|\cdot|_{H^{r}(\mathcal{O})}$. We use $(\cdot,\cdot)_{D}$ to represent the standard $L^{2}$-inner product over $D\subset\mathbb{R}^{d}$ and when $D$ coincides with $\Omega$, the subscript will be omitted. The corresponding norm is denoted as $\|\cdot\|_{L^{2}(D)}$. We define $\bm{H}(\mathrm{div};\Omega):=\{\bm{v}\in\bm{L}^{2}(\Omega),\nabla\cdot\bm{v}\in L^{2}(\Omega)\}$, which is equipped with the norm $\|\bm{v}\|_{\mathrm{div}}^{2}=\|\bm{v}\|_{L^{2}(\Omega)}^{2}+\|\nabla\cdot\bm{v}\|_{L^{2}(\Omega)}^{2}$. In addition, we use $\underline{H}(\mathrm{div};\Omega)$ to represent the tensor field where each row belongs to $\bm{H}(\mathrm{div};\Omega)$.
@@ 3.1 Preliminaries
where $\bm{t}_{i}:=(t_{i}^{1},t_{i}^{2})^{T},i=1,\cdots,4$ represents the unit tangential vector of $e_{i}$. Here $\bm{a}\times\bm{b}:=a_{1}b_{2}-a_{2}b_{1}$ for $\bm{a}=(a_{1},a_{2})^{T}$ and $\bm{b}=(b_{1},b_{2})^{T}$.
@@ 3.2 The discrete formulation and statement of the main results
Let $(\underline{\sigma},\bm{u},\gamma)$ be the exact solution of (2.1)-(2.2) and let $(\underline{\sigma}_{h},\bm{u}_{h},\gamma_{h})\in\underline{\Sigma}_{h}\times\bm{U}_{h}\times\Gamma_{h}$ be the discrete solution of (3.2)-(3.4) obtained by either Method 1 or Method 2. Assume that $(\underline{\sigma},\bm{u})\in\underline{H}^{1}(\Omega)\times\bm{H}^{1}(\Omega)$, then there exists a positive constant $C$ independent of the mesh size such that
$\|\underline{\sigma}-\underline{\sigma}_{h}\|_{L^{2}(\Omega)}$ $\leq Ch\Big{(}\|\underline{\sigma}\|_{H^{1}(\Omega)}+\|\gamma\|_{H^{1}(\Omega)}\Big{)},$
$\|\bm{u}-\bm{u}_{h}\|_{L^{2}(\Omega)}+\|\gamma-\gamma_{h}\|_{L^{2}(\Omega)}$ $\leq Ch\Big{(}\|\underline{\sigma}\|_{H^{1}(\Omega)}+\|\gamma\|_{H^{1}(\Omega)}\Big{)}.$
In addition, if we assume $\mathrm{div}\underline{\sigma}\in\bm{H}^{1}(\Omega)$, then the following superconvergence holds
$\|\bm{Q}_{h}\bm{u}-\bm{u}_{h}\|_{L^{2}(\Omega)}\leq Ch^{2}\Big{(}\|\underline{\sigma}\|_{H^{1}(\Omega)}+\|\gamma\|_{H^{1}(\Omega)}+\|\mathrm{div}\underline{\sigma}\|_{H^{1}(\Omega)}\Big{)},$
@@ 4 Error analysis
In this section, we show the unique solvability of the discrete formulation (cf. Theorem 4.1) and prove the convergence error estimates for $L^{2}$-error of stress and rotation; see Theorem 4.2.
@@ 4.1 Some elementary properties
$\|\bm{v}\|_{L^{2}(M)}$ and $\|\widehat{\bm{v}}\|_{L^{2}(M)}$ are equivalent uniformly in $h$. Moreover, it is shown in [16, Lemma 5.5] that
@@
$(\bm{\mathcal{J}}\bm{v}\cdot\bm{n},\mu)_{e}=(\bm{v}\cdot\bm{n},\mu)_{e}\quad\forall\mu\in P_{0}(e),e\in\mathcal{F}_{pr}^{\frac{1}{2}}\cap\partial E,$ (4.2)
@@
$\|\bm{v}-\bm{\mathcal{J}}\bm{v}\|_{L^{2}(M)}\leq Ch|\bm{v}|_{H^{1}(M)}\quad\forall\bm{v}\in\bm{H}^{1}(M),M\in\mathcal{T}_{M}.$
@@
where $e_{k}\in\mathcal{F}_{pr}^{\frac{1}{2}}$ is the half edge of $e$. In addition, the following holds (cf. [28, (3.9)])
@@
For a vector field $\bm{v}:=(v_{1},v_{2})^{T}$, we define
$\mathrm{curl}\bm{v}=\begin{pmatrix}\partial_{y}v_{1}&-\partial_{x}v_{1}\\ \partial_{y}v_{2}&-\partial_{x}v_{2}\end{pmatrix}.$
We use $\underline{I}_{h}^{e}$ and $\underline{\Pi}_{h}$ to represent the extension of $\bm{\mathcal{J}}$ and $\bm{\Pi}_{h}$ to the matrix field, i.e., $\underline{I}_{h}^{e}=[\bm{\Pi}^{u}]^{2}$ and $\underline{\Pi}_{h}=[\bm{\Pi}_{h}]^{2}$. We can deduce from the definition of $\underline{\Pi}_{h}$ that
@@
Now we recall the following stabilized $\bm{Q}_{1}-Q_{0}$ pair for the Stokes equations (cf. [22]).
@@
We define the finite element spaces for each macro-element $M,M\in\mathcal{T}_{pr}$ as follows:
$\bm{V}_{h}(M):$ =$=\{\bm{q}=(q_{1},q_{2})^{T}\in\bm{H}_{0}^{1}(M):\bm{q}_{|E}\in\bm{Q}_{1}(E),\forall E\in\mathcal{T}_{h}\;\textnormal{such that}\;E\subset M\},$
$W_{h}(M):$ =$=\{w\in L^{2}(\Omega):w_{|E}\in Q_{0}(E),\forall E\in\mathcal{T}_{h}\;\textnormal{such that}\;E\subset M\},$
Then, there exists a unique solution $(\bm{\rho}_{h},p_{h}^{*})\in\bm{V}_{h}(M)\times W_{h}(M)$ to the following system
$(\nabla\bm{\rho}_{h},\nabla\bm{v})_{M}+(p_{h}^{*},\nabla\cdot\bm{v})_{M}$ $=(\bm{f},\bm{v})_{M}\quad\forall\bm{v}\in\bm{V}_{h}(M),$ (4.10)
$(\mathrm{div}\bm{\rho}_{h},q)_{M}+\beta\sum_{e\in\mathcal{F}_{dl}\cap M}h_{e}(\left\llbracket p_{h}^{*}\right\rrbracket,\left\llbracket q\right\rrbracket)_{e}$ $=(g,q)_{M}\quad\forall q\in W_{h}(M),$ (4.11)
where $\beta>0$ is the stabilization parameter. Moreover, it holds
$\|\bm{\rho}_{h}\|_{H^{1}(\Omega)}\leq C\|g\|_{L^{2}(\Omega)},$ (4.12)
where $\bm{\rho}_{h}$ over each macro-element is defined by (4.10)-(4.11).
@@
(inf-sup condition). For Method 2, there exists a positive constant $C$ independent of the mesh size such that
$\sup_{\underline{0}\neq\underline{w}_{h}\in\underline{\Sigma}_{h}}\frac{\sum_{e\in\mathcal{F}_{pr}^{\frac{1}{2}}}(\underline{w}_{h}\bm{n},\left\llbracket\bm{v}_{h}\right\rrbracket)_{e}+(\textnormal{as}(\underline{w}_{h}),\eta_{h})}{\|\underline{w}_{h}\|_{L^{2}(\Omega)}}\geq C\|(\bm{v}_{h},\eta_{h})\|_{h}\quad\forall(\bm{v}_{h},\eta_{h})\in\bm{U}_{h}\times\Gamma_{h}.$ (4.13)
@@
We define $\underline{\psi}_{2}\in\underline{\Sigma}_{h}$ by
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\underline{\psi}_{2}=\mathrm{curl}(\bm{\rho}_{h}).}$ (4.16)
Since $(\bm{\rho}_{h})_{|M}\in\bm{V}_{h}(M)$, it is easy to check that $\mathrm{curl}(\bm{\rho}_{h})\in\underline{\Sigma}_{h}$. Therefore, in view of (4.11) with $g=\eta_{h}-\text{as}(\underline{\psi}_{1})$, we can obtain
@@
Here we use the fact that $\eta_{h}$ is a constant function over each macro-element $M$, and thus, the jump term for pressure vanishes.
It follows from (4.12) and (4.15) that